gary is six years older than peter.
the sum of there ages 34.
What are there ages?

Answer by Alan3354(69443)

gary is six years older than peter.
the sum of there ages 34.
What are there ages?
---------------
G = P + 6
---
G - P = 6
G + P = 34
------------- Add